40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 1 21:19:51 2019, gautam, Update, LSC, PRMI locked on carrier PRMIlocked.png
    Reply  Wed Oct 2 11:52:16 2019, gautam, Update, LSC, PRMI Oplev loop checkout PRM_OLTF.pdf
Message ID: 14920     Entry time: Tue Oct 1 21:19:51 2019     Reply to this: 14924
Author: gautam 
Type: Update 
Category: LSC 
Subject: PRMI locked on carrier 

Summary:

The PRMI was locked with the carrier field resonant in the PRC 🙌. The lock is pretty stable (I only let it stay locked for ~10mins and then deliberately unlocked to see if I could readily re-lock, but it has stayed locked for the last ~20mins while I typed this up). See Attachment #1 for the DC power monitor StripTool for a short section of lock.

Details:

  • This is the opposite of the config we'd want usually for locking the IFO, but it is a useful configuration for setting the alignment of the vertex optics, and also to train angular feedforward filters, so I decided to try it out.
  • Some patient alignment work was required. I started with the single arm locks, maximized TRX/TRY with ASS, and then misaligned the ETMs and brought the PRM into alignment.
  • The PRM Oplev spot was roughly centerd on its QPD once I judged I was getting decent PRMI cavity flashes on the POP camera. The PRMI Oplev servo needs some tuning, it is currently susceptible to oscillations in Pitch.
  • The error signals used were: REFL11_I ---> PRCL and AS55_Q ---> MICH.
  • The whitening gains were: REFL11 --> +18 dB, AS55 ---> +6 dB.
  • Triggering was done using POPDC, this worked better for me than any of the RF signals (e.g. POP22/POP110). Trigger ON --> 200cts, Trigger OFF --> 100 cts.
  • The DCPD whitening gains may not be set correctly - I think I remember POPDC being ~4000 cts in this configuration, but it may also be that we are not well centered on the POP photodiode.
  • The dominant cause of the POP circulating power seems to be the usual angular instability ascribed to the TTs. The OAF model wasn't running tonight (and I didn't want to try starting it and have to do a full vertex FE reboot tonight) so I didn't get a chance to engage the angular FF.

Next (for LSC activities):

  • PRMI locking with the sidebands resonant in the PRC.
  • DRMI locking

I'm leaving the LSC mode off for tonight, but with the PRMI optics aligned and ETMs misaligned.

Attachment 1: PRMIlocked.png  128 kB  Uploaded Tue Oct 1 22:43:04 2019  | Hide | Hide all
PRMIlocked.png
ELOG V3.1.3-